# RGFN: Synthesizable Molecular Generation Using GFlowNets

### v1

@@ 4.4 Examination of the produced ligands
In the final stage of experiments we examine the capabilities of RGFN to produce high quality ligands across multiple diverse docking targets (see Appendix F for more details). The aim is to evaluate whether 1) the chemical language used is expressive enough to produce structurally diverse molecules for different targets, and 2) whether the generated ligands form realistic poses in the binding pockets. We first demonstrate the diversity of ligands across targets on a UMAP plot of extended-connectivity fingerprints (Figure 6). Ligands assigned to specific targets form very distinct clusters, showcasing their diversity. Interestingly, we also observe structural differences between sEH proxy and sEH docking, possibly indicating poor approximation of docking scores by the proxy model. Secondly, we examine the docking poses of the highest scoring generated ligands (Figure 7). As can be seen, the generated molecules produce realistic docking poses, closely resembling the poses of known ligands (Appendix J), despite being diverse in terms of structural similarity (Appendix L). We further conduct a cost analysis and synthesis planning for top modes in Appendices M and N. Overall, this demonstrates the usefulness of the proposed RGFN approach in the docking-based screens.
Figure 6: UMAP plot of chemical structures of top-500 modes generated for each target. RGFN generates sufficient chemical diversity to produce distinct clusters of compounds. See Appendix F for description of each target protein.
@@ 5 Limitations
The current set of building blocks and reactions tends to generate linear and flat-shaped molecules. Adding a small set of cyclization reactions (such as peptide macrocyclization and ring-closing metathesis), along with more complex-shaped scaffold building blocks, as well as reactions that introduce sp3 hybridized atoms and stereochemical complexity will therefore allow for greater shape diversity and the generation of more potent molecules [44, 17]. It is also important to recognize that RGFN does not explicitly generate synthetic routes to the molecules.
Another significant limitation in the quality of the generated molecular structures is the reliance on molecular docking as a scoring oracle. Although molecular docking has been successfully used in large-scale virtual screening efforts [38, 52, 31], it has well-known shortcomings in its predictive power. First, docking scores correlate strongly with molecular weight (MW) [12] and do not account for drug-likeness requirements like optimal MW or ClogP. In this work, molecule size was constrained only by the number of reaction steps, encouraging RGFN to generate large molecules within the building block limit. This can be somewhat rectified by augmenting reward with a drug-likeness or ligand efficiency term. Second, its binding affinity predictions and rankings often correlate weakly with experimental values and are highly dependent on the nature of the target protein’s binding site [63, 61]. This is further illustrated by the fact that known ligands are not necessarily characterized by highest possible docking scores (Appendix K). This limitation impacts the learning of the chemical structure-activity relationship space, leading to the generation of sub-optimal molecules. One solution to this limitation is to incorporate more accurate but computationally expensive methods (such as ensemble docking, MM-PBSA, and FEP) within a multi-fidelity framework [23]. However, since we focus on robust, affordable, and facile synthesis methods, we ultimately aim to extend our approach beyond computational scoring methods by directly conducting experimental evaluation of synthesized compound batches within an active learning loop.
@@ 6 Conclusions
This work was supported by funding from CQDM Fonds d’Accélération des Collaborations en Santé (FACS) / Acuité Québec and the National Research Council (NRC) Canada, the Canadian Institutes for Health Research (CIHR), Samsung and Microsoft. Computational resources were provided by the Digital Research Alliance of Canada (https://alliancecan.ca) and Mila (https://mila.quebec). The research of P. Gainski was supported by the National Science Centre (Poland), grant no. 2022/45/B/ST6/01117. We gratefully acknowledge Poland’s high-performance Infrastructure PLGrid (ACK Cyfronet Athena, HPC) for providing computer facilities and support within computational grant no PLG/2023/016550. Additional thanks for funding provided to the University of Toronto’s Acceleration Consortium from the Canada First Research Excellence Fund (Grant number - CFREF-2022-00042).
@@ B.2 Generative models
All sampling algorithms were outfitted with the Vina GPU-2.1 docking, senolytic proxy, and sEH proxy scoring functions. While model architecture hyperparameters and batch sizes were kept consistent between FGFN and RGFN, we allowed FGFN a maximum fragment count of 6 as opposed to RGFN’s 5 due to RGFN’s larger average building block sizes.
@@
For SyntheMol experiments, we used the default building block library of 132,479 compatible molecules and pre-computed docking, senolytic, and sEH proxy scores for all prior to executing rollouts to follow the established methodology. Due to CPU constraints, sampling 500,000 molecules with SyntheMol was impractical. Instead, we executed 100,000 rollouts over approximately 72 hours to match the RGFN training time with docking, yielding 111,964 unique molecules. Additionally, we performed 50,000 rollouts each (approximately 30 hours) for sEH and senolytic proxies, resulting in 73,941 and 69,652 unique molecules, respectively.
@@ C.2 Boron substitution for docking
Appendix D Implicit reactions
@@
Figure 8: An example of an implicit reaction: urea synthesis reaction encoded in SMARTS (reaction index 56).
In addition to urea formation (reaction indexes 56, 57, and 58), other implicit reactions encoded in SMARTS include azide-alkyne cycloaddition (indexes 44 and 46), azide-nitrile cycloadditions (indexes 45 and 47), and tetrazole synthesis using peptide terminal thiourea cyclizations (index 60).
Appendix E Sampling large fragment set from Enamine building blocks
@@
Appendix F Target selection
@@
Appendix G Ligand post-processing
To ensure the diversity, specificity, and conformer validity of top-generated molecules for each target, we initially categorized our molecules into distinct modes, each representing any SMILES string with a Tanimoto similarity of 0.5 or lower with all other modes. Subsequently, we selected the top 100 modes based on their Vina-GPU 2.1 scores and filtered their docked poses using PoseBusters [11] in "mol" mode, where any pose failing any PoseBusters check was excluded from consideration. As a final precaution, we selected only modes with Tanimoto coefficients to known aggregators of 0.4 or lower using UCSF’s Aggregation Advisor [25] dataset. This process resulted in 35, 68, 31, and 15 top modes for sEH, ClpP, Mpro, and TBLR1 binders, respectively Appendix I. Comparative analyses of docked top RGFN modes and confirmed sEH, ClpP, and Mpro ligand poses can be found in Appendices J, K and L.
Appendix H Posebusters, Aggregation Advisor analysis of top 100 modes
@@
Appendix I Top filtered molecules for all targets
@@
Appendix J Docked poses of top generated molecules
@@
Appendix K Docking analysis of known ligands
@@
Appendix L Tanimoto similarity to known ligands
@@
Appendix M Ligands cost analysis
To demonstrate the synthesizability and cost-effectiveness of the ligands produced by RGFN, we conducted a comprehensive cost analysis and proposed a synthesis plan for the top 10 scored ClpP hits generated by RGFN and SyntheMol (see Appendix N). We compared these two methods due to their similarity in chemical approach. For the cost calculation, we only considered the price of building blocks, which were directly sourced from EnamineStore for the case of Synthemol ligands. All prices are represented in US dollars per 0.1 mmol of product (Figures 19 and 20). Notably, for very common reagents only available in bulk, costs per gram were estimated by taking the cheapest or smallest available alternative and dividing its cost by its mass. All prices for individual building blocks were sourced from their respective vendors, which included Millipore Sigma, Oakwood Chemicals, Combi-Blocks, TCI, and AngeneSci.
@@
Appendix N Examples of plausible synthetic routes to molecules

### v2

@@ 4.4 Examination of the produced ligands
In the final stage of experiments we examine the capabilities of RGFN to produce high quality ligands across multiple diverse docking targets (see Appendix G for more details). The aim is to evaluate whether 1) the chemical language used is expressive enough to produce structurally diverse molecules for different targets, and 2) whether the generated ligands form realistic poses in the binding pockets. We first demonstrate the diversity of ligands across targets on a UMAP plot of extended-connectivity fingerprints (Figure 6). Ligands assigned to specific targets form very distinct clusters, showcasing their diversity. Interestingly, we also observe structural differences between sEH proxy and sEH docking, possibly indicating poor approximation of docking scores by the proxy model. Secondly, we examine the docking poses of the highest scoring generated ligands (Figure 7). As can be seen, the generated molecules produce realistic docking poses, closely resembling the poses of known ligands (Appendix K), despite being diverse in terms of structural similarity (Appendix M). We further conduct a cost analysis and synthesis planning for top modes in Appendices N and O. Overall, this demonstrates the usefulness of the proposed RGFN approach in the docking-based screens.
Figure 6: UMAP plot of chemical structures of top-500 modes generated for each target. RGFN generates sufficient chemical diversity to produce distinct clusters of compounds. See Appendix G for description of each target protein.
@@ 5 Limitations
The current set of building blocks and reactions tends to generate linear and flat-shaped molecules. Adding a small set of cyclization reactions (such as peptide macrocyclization and ring-closing metathesis), along with more complex-shaped scaffold building blocks, as well as reactions that introduce sp3 hybridized atoms and stereochemical complexity will therefore allow for greater shape diversity and the generation of more potent molecules [51, 20]. It is also important to recognize that RGFN does not explicitly generate synthetic routes to the molecules in a strict chemical sense, which in addition to a sequence of reactions transforming sets of reactants into products (which RGFN does provide), would also include reaction conditions, external reagents, catalysts, protection group strategies, and other important considerations.
Another significant limitation in the quality of the generated molecular structures is the reliance on molecular docking as a scoring oracle. Although molecular docking has been successfully employed in large-scale virtual screening efforts [44, 61, 35], it has well-known shortcomings in its predictive power. First, docking scores correlate strongly with molecular weight (MW) [13] and do not account for drug-likeness requirements like optimal MW or ClogP. In this work, molecule size was constrained only by the number of reaction steps, encouraging RGFN to generate large molecules within the building block limit. This can be somewhat rectified by augmenting reward with a drug-likeness or ligand efficiency term. Second, its binding affinity predictions and rankings often correlate weakly with experimental values and are highly dependent on the nature of the target protein’s binding site [73, 71]. This is further illustrated by the fact that known ligands are not necessarily characterized by highest possible docking scores (Appendix L). This limitation impacts the learning of the chemical structure-activity relationship space, leading to the generation of sub-optimal molecules. One solution to this limitation is to incorporate more accurate but computationally expensive methods (such as ensemble docking, MM-PBSA, and FEP) within a multi-fidelity framework [26]. However, since we focus on robust, affordable, and facile synthesis methods, we ultimately aim to extend our approach beyond computational scoring methods by directly conducting experimental evaluation of synthesized compound batches within an active learning loop.
@@ 6 Conclusions
This work was supported by funding from CQDM Fonds d’Accélération des Collaborations en Santé (FACS) / Acuité Québec and the National Research Council (NRC) Canada, the Canadian Institutes for Health Research (CIHR), grant no. FDN-167277, Samsung and Microsoft. D. Shevchuk is grateful for support from the Mitacs Globalink Graduate Fellowship, grant no. FR121160/FR121161. The research of P. Gaiński was supported by the National Science Centre (Poland), grant no. 2022/45/B/ST6/01117. Additional thanks for funding provided to the University of Toronto’s Acceleration Consortium from the Canada First Research Excellence Fund, grant no. CFREF-2022-00042. Computational resources were provided by the Digital Research Alliance of Canada (https://alliancecan.ca) and Mila (https://mila.quebec). We gratefully acknowledge Poland’s high-performance Infrastructure PLGrid (ACK Cyfronet Athena, HPC) for providing computer facilities and support within computational grant no PLG/2023/016550.
@@ B.1 Proxy models
The DRD2 proxy is described in [53, 28]. It is a support vector machine classifier with a Gaussian kernel using ECFP6 fingerprints as a feature representation.
@@ B.2 Generative models
For FGFN+SA, we used a modified reward function $R(x)=exp(\beta*(0.5*proxy(x)/max\_proxy+(10-SA\_score)/10)$, with $\beta$ adjusted per proxy to match the original reward range.
All sampling algorithms were outfitted with the Vina GPU-2.1 docking, senolytic proxy, sEH proxy, and DRD2 proxy scoring functions. While model architecture hyperparameters and batch sizes were kept consistent between FGFN and RGFN, we allowed FGFN a maximum fragment count of 6 as opposed to RGFN’s 5 due to RGFN’s larger average building block sizes.
@@
For SyntheMol experiments, we used the default building block library of 132,479 compatible molecules and pre-computed docking, senolytic, and sEH proxy scores for all prior to executing rollouts to follow the established methodology. Due to CPU constraints, sampling 500,000 molecules with SyntheMol was impractical. Instead, we executed 100,000 rollouts over approximately 72 hours to match the RGFN training time with docking, yielding 111,964 unique molecules. Additionally, we performed 50,000 rollouts each (approximately 24 hours) for sEH, senolytic, and DRD2 proxies, resulting in 73,941, 69,652, and 62,320 unique molecules, respectively.
casVAE was trained with Bayesian optimization (BO) using default parameters consisting of a hidden size of 200, latent size of 50, and message passing depth of 2. Again, due to time constraints imposed by BO latent space updates, we instead opted to approximately match RGFN training times, training for 72, 24, 24, and 24 hours on the docking, sEH, senolytic, and DRD2 tasks respectively for a total of 135, 41, 38, and 40 rollouts and 7708, 2097, 1521, and 2165 total generated molecules, respectively.
@@ C.2 Boron substitution for docking
Appendix D Computational resources used
Evaluating fragment scaling took approximately 800 GPU hours on GeForce RTX 4090 in total. Remaining experiments took roughly 24 GPU hours per run on Quadro RTX 8000 for sEH, DRD2 and senolytic proxies, and roughly 72 GPU hours per run on an A100 for docking-based proxies.
Appendix E Implicit reactions
@@
Figure 8: An example of an implicit reaction: urea synthesis reaction encoded in SMARTS.
In addition to urea formation, other implicit reactions encoded in SMARTS include azide-alkyne cycloaddition, azide-nitrile cycloadditions, and tetrazole synthesis using peptide terminal thiourea cyclizations.
Appendix F Sampling large fragment set from Enamine building blocks
@@
Appendix G Target selection
@@
Appendix H Ligand post-processing
To ensure the diversity, specificity, and conformer validity of top-generated molecules for each target, we initially categorized our molecules into distinct modes, each representing any SMILES string with a Tanimoto similarity of 0.5 or lower with all other modes. Subsequently, we selected the top 100 modes based on their Vina-GPU 2.1 scores and filtered their docked poses using PoseBusters [11] in "mol" mode, where any pose failing any PoseBusters check was excluded from consideration. As a final precaution, we selected only modes with Tanimoto coefficients to known aggregators of 0.4 or lower using UCSF’s Aggregation Advisor [29] dataset. This process resulted in 35, 68, 31, and 15 top modes for sEH, ClpP, Mpro, and TBLR1 binders, respectively Appendix J. Comparative analyses of docked top RGFN modes and confirmed sEH, ClpP, and Mpro ligand poses can be found in Appendices K, L and M.
Appendix I Posebusters, Aggregation Advisor analysis of top 100 modes
@@
Appendix J Top filtered molecules for all targets
@@
Appendix K Docked poses of top generated molecules
@@
Appendix L Docking analysis of known ligands
@@
Appendix M Tanimoto similarity to known ligands
@@
Appendix N Ligands cost analysis
To demonstrate the synthesizability and cost-effectiveness of the ligands produced by RGFN, we conducted a comprehensive cost analysis and proposed a synthesis plan for the top 10 scored ClpP hits generated by RGFN and SyntheMol (see Appendix O). We compared these two methods due to their similarity in chemical approach. For the cost calculation, we only considered the price of building blocks, which were directly sourced from EnamineStore for the case of Synthemol ligands. All prices are represented in US dollars per 0.1 mmol of product (Figures 19 and 20). Notably, for very common reagents only available in bulk, costs per gram were estimated by taking the cheapest or smallest available alternative and dividing its cost by its mass. All prices for individual building blocks were sourced from their respective vendors, which included Millipore Sigma, Oakwood Chemicals, Combi-Blocks, TCI, and AngeneSci.
@@
Appendix O Examples of plausible synthetic routes to molecules